# Double-coset zeta functions for groups acting on trees

### v1

@@ 5.4. Explicit formulae: the weakly locally $\infty$-transitive case
Let $(G,T)$ be a weakly locally $\infty$-transitive group action on a locally finite tree with quotient map $\pi\colon T\to\Gamma=G\backslash T$. Assume that $\Gamma$ is finite, and that its standard edge weight $\omega$ takes values in $\mathbb{Z}_{\geq 2}$ and satisfies $\omega(a)\geq 3$ or $\omega(\bar{a})\geq 3$ for every $a\in E\Gamma$. Let also $N_{\mathrm{edg}}=N_{\mathrm{edg}}^{\omega}$ and $N_{\mathrm{vert}}=N_{\mathrm{vert}}^{\omega}$ be as in Definition 4.4.
@@
    We implicitly set a total order on $E\Gamma$. Thus, provided $|E\Gamma|<\infty$, we can regard $\mathcal{E}(s)$ as a $|E\Gamma|$-dimensional matrix $[\mathcal{E}(s)(a,b)]_{a,b\in E\Gamma}$ with complex entries, and the $e_{u}$’s in (5.13) as row vectors in $\mathbb{C}^{|E\Gamma|}$. For all $a,b\in E\Gamma$, note that $e_{a}\mathcal{E}(s)e_{b}^{t}=\mathcal{E}(s)(a,b)$.
The term “Bass operator” is taken after [9, Definition 3.10]. The reader is referred to Section 7.3 for further connections with [9].
@@
Let $\mathcal{E}(s)$ be as in Definition 5.10. For every $n\geq 1$, let $\mathcal{E}(s)^{n}$ be the $n$-th power of $\mathcal{E}(s)$, and $\mathcal{E}(s)^{0}$ be the identity operator on $\mathbb{C}\llbracket E\Gamma\rrbracket$. Then, for all $n\geq 0$ and $a,b\in E\Gamma$, we observe that
$e_{a}\mathcal{E}(s)^{n}e_{b}^{t}=\mathcal{E}(s)^{n}(a,b)=\sum_{\mathfrak{p}\in\mathcal{P}_{\Gamma}(a\to b)\atop\ell(\mathfrak{p})=n+1}N_{\mathrm{edg}}(\mathfrak{p})^{-s}.$
@@
Setting [$\Gamma$].
Let $\Gamma$ be a finite connected non-empty graph with an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 2}$ satisfying $\omega(a)\geq 3$ or $\omega(\bar{a})\geq 3$, for every $a\in E\Gamma$.
@@
Let $(\Gamma,\omega)$ be an edge-weighted graph satisfying Setting [$\Gamma$]. Then, for all $u,w\in\Gamma$,
@@
where $I$ be the identity matrix in $\mathrm{Mat}_{|E\Gamma|}(\mathbb{C})$ and
@@
In particular, for all $u,w\in\Gamma$ the function $\mathcal{Z}_{\Gamma,u\to w}(s)$ is a meromorphic function over $\mathbb{C}$.
@@
Fact 5.13 (Matrix Determinant Lemma, cf. [15]).
@@
Let $s\in\mathbb{C}$ be such that $\sum_{n=0}^{\infty}\mathcal{E}(s)$ converges. Recall that $\sum_{n=0}^{\infty}\mathcal{E}(s)=(I-\mathcal{E}(s))^{-1}$. By Remark 5.11.1, if $u\in V\Gamma$ then
$\begin{split}&\mathcal{Z}_{\Gamma,u\to w}(s)=\\ &=\left\{\begin{array}[]{ll}\displaystyle{\mathbbm{1}_{\{u\}}(w)+\sum_{n=0}^{\infty}\sum_{a\in o^{-1}(u)}\omega(a)^{-s}e_{a}\mathcal{E}(s)^{n}e_{w}^{t}},&\text{if }w\in V\Gamma;\\ \displaystyle{\mathbbm{1}_{o^{-1}(u)}(w)\cdot\omega(w)^{-s}+\sum_{n=1}^{\infty}\sum_{a\in o^{-1}(u)}\omega(a)^{-s}e_{a}\mathcal{E}(s)^{n}(e_{w}+e_{\bar{w}})^{t}},&\text{if }w\in E\Gamma.\end{array}\right.\end{split}$
@@
$\begin{split}\mathcal{Z}_{\Gamma,u\to w}(s)=\left\{\begin{array}[]{ll}\displaystyle{\mathbbm{1}_{\{o(u),t(u)\}}(w)+\sum_{n=1}^{\infty}(e_{u}+e_{\bar{u}})\mathcal{E}(s)^{n}e_{w}^{t}},&\text{if }w\in V\Gamma;\\ \displaystyle{\mathbbm{1}_{\{u,\bar{u}\}}(w)+\sum_{n=1}^{\infty}(e_{u}+e_{\bar{u}})\mathcal{E}(s)^{n}(e_{w}^{t}+e_{\bar{w}}^{t})},&\text{if }w\in E\Gamma.\end{array}\right.\end{split}$
@@
Let $\Gamma$ be a $1$-segment with $E\Gamma=\{a,\bar{a}\}$, $c=o(a)$ and $d=t(a)$. Set $\omega(a):=\alpha+1$ and $\omega(\bar{a}):=\beta+1$, where $\alpha,\beta\in\mathbb{Z}_{\geq 1}$ with $\alpha\geq 2$ or $\beta\geq 2$. Set the order $\leq$ on $E\Gamma$ such that $a<\bar{a}$, and identify $\mathbb{C}\llbracket E\Gamma\rrbracket$ with $\mathbb{C}^{2}$, $e_{a}=e_{d}$ with $(1,0)$ and $e_{\bar{a}}=e_{c}=(0,1)$. Then, for every $s\in\mathbb{C}$,
$\begin{split}\mathcal{E}(s)&=\begin{bmatrix}\mathcal{E}(s)(a,a)&\mathcal{E}(s)(a,\bar{a})\\ \mathcal{E}(s)(\bar{a},a)&\mathcal{E}(s)(\bar{a},\bar{a})\end{bmatrix}=\begin{bmatrix}0&\beta^{-s}\\ \alpha^{-s}&0\end{bmatrix};\\ \mathcal{U}_{c,c}(s)&=(\alpha+1)^{-s}e_{c}^{t}\cdot e_{a}=\begin{bmatrix}0&0\\ (\alpha+1)^{-s}&0\end{bmatrix};\\ \mathcal{U}_{a,a}(s)&=(e_{a}+e_{\bar{a}})^{t}\cdot(e_{a}+e_{\bar{a}})\cdot\mathcal{E}(s)=\begin{bmatrix}\alpha^{-s}&\beta^{-s}\\ \alpha^{-s}&\beta^{-s}\end{bmatrix}.\end{split}$
@@
In particular, let $(G,T)$ be a locally $\infty$-transitive action on a locally finite tree with quotient graph $\Gamma$ and standard edge weight $\omega$ as before. By Remark 5.9.2 and (5.17), we have explicit formulae for $\zeta_{G,G_{v},G_{v}}(s)$ and $\zeta_{G,G_{e},G_{e}}(s)$ for all $v\in VT$ with $G\cdot v=c$ and $e\in ET$ with $G\cdot e=a$. For instance, one may take $G=\mathrm{SL}_{2}(\mathbb{Q}_{p})$ and $T$ the Bruhat–Tits tree of $G$ (cf. Example 3.3(ii)). In this case $\alpha=\beta=p$. Let $v\in VT$ be the vertex with $G_{v}=\mathrm{SL}_{2}(\mathbb{Z}_{p})$, and $e\in ET$ be the edge whose pointwise stabiliser is the standard Iwahori subgroup. Then,
@@
Other examples can be obtained from Example 3.12. Note that the formulae before agree with [8, Example 1.7] in case that $G$ is the group of automorphisms of a bi-coloured tree $T$.
@@
$\begin{split}\mathcal{E}(s)&=\begin{bmatrix}\mathcal{E}(s)(a,a)&\mathcal{E}(s)(a,\bar{a})\\ \mathcal{E}(s)(\bar{a},a)&\mathcal{E}(s)(\bar{a},\bar{a})\end{bmatrix}=\begin{bmatrix}(\alpha+1)^{-s}&\beta^{-s}\\ \alpha^{-s}&(\beta+1)^{-s}\end{bmatrix};\\ \mathcal{U}_{c,c}(s)&=(\alpha+1)^{-s}e_{c}^{t}\cdot e_{a}+(\beta+1)^{-s}e_{c}^{t}\cdot e_{\bar{a}}=\begin{bmatrix}(\alpha+1)^{-s}&(\beta+1)^{-s}\\ (\alpha+1)^{-s}&(\beta+1)^{-s}\end{bmatrix};\\ \mathcal{U}_{a,a}(s)&=(e_{a}+e_{\bar{a}})^{t}\cdot(e_{a}+e_{\bar{a}})\mathcal{E}(s)=\begin{bmatrix}(\alpha+1)^{-s}+\alpha^{-s}&(\beta+1)^{-s}+\beta^{-s}\\ (\alpha+1)^{-s}+\alpha^{-s}&(\beta+1)^{-s}+\beta^{-s}\end{bmatrix}.\end{split}$
@@ 5.5. Explicit formulae: the (P)-closed case
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram based on a non-empty finite connected graph $\Gamma$. Choose an inversion $\iota$ in $\Delta$ and $c_{0}\in V\Gamma$. Denote by $(T=T(\Delta,\iota,c_{0}),\pi,\mathcal{L})$ the standard $\Delta$-tree associated to $\iota$ and $c_{0}$, let $v_{0}$ be the root of $T$ (cf. Section 3.4.1), and set $G=U(\Delta,\iota,c_{0})$. Assume that the standard edge weights on $\Gamma$ and $\Delta$, denoted by $\omega$ and $\mathcal{W}$ respectively, take values in $\mathbb{Z}_{\geq 2}$ and $\mathbb{Z}_{\geq 0}$, respectively. Finally, assume that $(G,T)$ has property $(\ast_{k})$ for some $k\geq 1$.
Setting [(P)-cl] guarantees that the series defining $\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)$ converges at some $s\in\mathbb{C}$, for all $t_{1},t_{2}\in T$ (cf. Theorem 5.7). Thanks to the following remark, from now on we may focus only on the case in which $t_{1}\in\{v_{0}\}\cup o^{-1}(v_{0})$ while studying $\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)$.
Remark 5.15.1.
Let $\Delta$ be a local action diagram and $\mathbb{T}=(T,\pi,\mathcal{L})$ be $\Delta$-tree. Consider an inversion $\iota$ in $\Delta$ and denote by $(T(\Delta,\iota,c_{0}),\pi_{0},\mathcal{L}_{0})$ the standard $\Delta$-tree associated to $\iota$ and some $c_{0}\in V\Gamma$. By Theorem 3.8(i), there is a graph isomorphism $\phi\colon T\to T(\Delta,\iota,c_{0})$ such that $\pi_{0}=\pi\circ\phi$ and $U(\Delta,\iota,c_{0})=\phi U(\Delta,\mathbb{T})\phi^{-1}$. Set $G=U(\Delta,\mathbb{T})$ and $H=U(\Delta,\iota,c_{0})$. Then, for all $t\in T$ the following map is bijective:
$|G_{t_{1}}\cap G_{t_{1}}\cap gG_{t_{2}}g^{-1}|=|H_{\phi(t_{1})}:H_{\phi(t_{1})}\cap hH_{\phi(t_{2})}h^{-1}|$
and then
$\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)=\zeta_{H,H_{\phi(t_{1})},H_{\phi(t_{2})}}(s),$
whenever one series before is defined.
$\zeta_{G,G_{v},G_{t}}(s)=\zeta_{H,H_{v_{0}},H_{\phi(t)}}(s).$
Moreover, for all $e\in o^{-1}(v)$ we have $\phi(e)\in o^{-1}(v_{0})$ and, for all $t\in T$,
$\zeta_{G,G_{e},G_{t}}(s)=\zeta_{H,H_{\phi(e)},H_{\phi(t)}}(s).$
@@
(5.20) $\zeta_{G,G_{v_{0}},G_{t}}(s)=\left\{\begin{array}[]{ll}\displaystyle{\mathbbm{1}_{\{c_{0}\}}(u)+\sum_{a\in o^{-1}(c_{0}),\atop\xi\in\mathcal{P}_{(\Delta,\iota)}(X_{a}\to X_{U})}\omega(a)^{-s-1}\mathcal{W}(\xi)^{-s-1},}&\text{if }u\in V\Gamma;\\ \displaystyle{\sum_{a\in o^{-1}(c_{0}),\atop\xi\in\mathcal{P}_{(\Delta,\iota)}(X_{a}\to X_{U})}\omega(a)^{-s-1}\mathcal{W}(\xi)^{-s-1},}&\text{if }u\in E\Gamma.\end{array}\right.$
Moreover, for every $e\in ET$ with $\pi(e)=a$ and $\mathcal{L}(e)=x\in X_{c_{0}}$, we have
(5.21) $\zeta_{G,G_{e},G_{t}}(s)=\eta_{a}(u)+\sum_{\xi\in\mathcal{P}_{(\Delta,\iota)}(x\to X_{U}),\atop\ell(\xi)\geq 2}\mathcal{W}(\xi)^{-s-1}+\sum_{y\in X_{c_{0}}\setminus\{x\},\atop\xi\in\mathcal{P}_{(\Delta,\iota)}(y\to X_{U}),\,\ell(\xi)\geq 1}\mathcal{W}_{\mathrm{rev}}(x,y)^{-s-1}\mathcal{W}(\xi)^{-s-1},$
@@
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram together with a function $\mathcal{W}\colon X\times X\to\mathbb{Z}_{\geq 0}$ (recall that $X=\bigsqcup_{a\in E\Gamma}X_{a}$). Let $\mathbb{C}\llbracket X\rrbracket$ be the complex vector space of all formal sums $\sum_{x\in X}\gamma_{x}x$, where $\gamma_{x}\in\mathbb{C}$ for every $x\in X$. For every non-empty set $S\subseteq X$, let
$f_{S}=\sum_{x\in S}x\in\mathbb{C}\llbracket X\rrbracket.$
@@
In what follows, we implicitly fix a total order on $X$. Thus, we regard $\mathcal{F}(s)$ and the $f_{x}$’s as a $|X|$-dimensional matrix $[\mathcal{F}(s)(x,y)]_{x,y\in X}$ and as $|X|$-dimensional row vectors with complex entries, respectively. For all $x,y\in X$, note that $f_{x}\mathcal{F}(s)f_{y}^{t}=\mathcal{F}(s)(x,y)$.
@@
Let $\mathcal{F}(s)$ be as in Definition 5.17. For every $n\geq 1$, let $\mathcal{F}(s)^{n}$ be the $n$-th power of $\mathcal{F}(s)$, and $\mathcal{F}(s)^{0}$ be the identity operator on $\mathbb{C}\llbracket X\rrbracket$. For $n\geq 0$ and for all $x,y\in X$, we claim that
(5.22) $f_{x}\cdot\mathcal{F}(s)^{n}\cdot f_{y}^{t}=\mathcal{F}(s)^{n}(x,y)=\sum_{\xi\in\mathcal{P}_{(\Delta,\iota)}(x\to y):\,\ell(\xi)=n+1}\mathcal{W}(\xi)^{-s},$
@@
Let $G=U(\Delta,\iota,c_{0})$ and $T=T(\Delta,\iota,c_{0})$ be as in Setting [(P)-cl]. Let $t\in T$ with $\pi(t)=u$ and $e\in o^{-1}(v_{0})$ with $\mathcal{L}(e)=x$. Then, for every $r\in\{v_{0},e\}$, we have
@@
$\kappa_{\pi(r)}(u)=\left\{\begin{array}[]{cl}\mathbbm{1}_{\{c_{0}\}}(u)-1,&\text{if }r=v_{0};\\ \mathbbm{1}_{X_{A}}(x)-1,&\text{if }r=e.\end{array}\right.$
@@
One proceeds analogously as in the proof of Theorem 5.12. Let $s\in\mathbb{C}$ such that $\sum_{n=0}^{\infty}\mathcal{F}(s+1)$ converges. By Proposition 5.16 and Remark 5.18.1, we deduce what follows:
$\begin{split}&\zeta_{G,G_{v},G_{t}}(s)=\\ &=\left\{\begin{array}[]{ll}\displaystyle{\mathbbm{1}_{\{c_{0}\}}(u)+\sum_{a\in o^{-1}(c_{0})}\omega(a)^{-s-1}f_{X_{a}}\Bigg{(}\sum_{n=0}^{\infty}\mathcal{F}(s+1)^{n}\Bigg{)}f_{X_{U}}^{t},}&\text{if }u\in V\Gamma;\\ \displaystyle{\sum_{a\in o^{-1}(c_{0})}\omega(a)^{-s-1}f_{X_{a}}\Bigg{(}\sum_{n=0}^{\infty}\mathcal{F}(s+1)^{n}\Bigg{)}f_{X_{U}}^{t}},&\text{if }u\in E\Gamma;\end{array}\right.\\ &\zeta_{G,G_{e},G_{t}}(s)=\eta_{a}(u)+f_{x}\Bigg{(}\sum_{n=1}^{\infty}\mathcal{F}(s+1)^{n}\Bigg{)}f_{X_{U}}^{t}+\\ &\hskip 62.59596pt+\sum_{y\in X_{c_{0}}\setminus\{x\}}\mathcal{W}_{\mathrm{rev}}(x,y)^{-s-1}f_{y}\Bigg{(}\sum_{n=0}^{\infty}\mathcal{F}(s+1)^{n}\Bigg{)}f_{X_{U}}^{t}=\\ &=\mathbbm{1}_{X_{U}}(x)+\Big{(}f_{x}\mathcal{F}(s+1)+\sum_{y\in X_{c_{0}}\setminus\{x\}}\mathcal{W}_{\mathrm{rev}}(x,y)^{-s-1}f_{y}\Big{)}\Bigg{(}\sum_{n=0}^{\infty}\mathcal{F}(s+1)^{n}\Bigg{)}f_{X_{U}}^{t}.\end{split}$
Since $\sum_{n=0}^{\infty}\mathcal{F}(s+1)^{n}=(I-\mathcal{F}(s+1))^{-1}$, Fact 5.13 yields the claim. ∎

### v2

@@ 5.4. Explicit formulae: the weakly locally $\infty$-transitive case
Let $(G,T)$ be a weakly locally $\infty$-transitive group action on a locally finite tree as in Convention 3.0.1 and with quotient map $\T\pi\colon T\to\Gamma=G\backslash$ T. Assume that $\Gamma$ is finite, and that its standard edge weight $\omega$ takes values in $\mathbb{Z}_{\geq 2}$ (so $T$ has no leaves) and satisfies $\omega(a)\geq 3$ or $\omega(\bar{a})\geq 3$ for every $a\in E\Gamma$. Let also $N_{\mathrm{edg}}=N_{\mathrm{edg}}^{\omega}$ and $N_{\mathrm{vert}}=N_{\mathrm{vert}}^{\omega}$ be as in Definition 4.4.
@@
    We implicitly set a total order on $E\Gamma$. Thus, provided $|E\Gamma|<\infty$, we can regard $\mathcal{E}(s)$ as a $|E\Gamma|$-dimensional matrix $\mathcal{E}(s)(a,b)]_{a,b\in E\Gamma}$ with complex entries, and the $e_{u}$’s in (5.11) as row vectors in $\mathbb{C}^{|E\Gamma|}$. For all $a,b\in E\Gamma$, note that $e_{a}\mathcal{E}(s)e_{b}^{t}=\mathcal{E}(s)(a,b)$.
The term “Bass operator” is taken after [11, Definition 3.10]. The reader is referred to Section 7.3 for further connections with [11].
@@
Let $\mathcal{E}(s)$ be as in Definition 5.12 and assume that $\Gamma$ is finite. For every $n\geq 1$, let $\mathcal{E}(s)^{n}$ be the $n$-th power of $\mathcal{E}(s)$, and $\mathcal{E}(s)^{0}$ be the identity operator on $\mathbb{C}\llbracket E\Gamma\rrbracket$. Then, for all $n\geq 0$ and $a,b\in E\Gamma$, we observe that
$\mathcal{E}(s)^{n}(a,b)=\sum_{\mathfrak{p}\in\mathcal{P}_{\Gamma}(a\to b)\atop\ell(\mathfrak{p})=n+1}N_{\mathrm{edg}}(\mathfrak{p})^{-s}.$
@@
Setting $\Gamma$].
Let $\Gamma$ be a finite connected graph with $E\Gamma\neq\emptyset$ and with an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 2}$ satisfying $\omega(a)\geq 3$ or $\omega(\bar{a})\geq 3$, for every $a\in E\Gamma$.
@@
Let $(\Gamma,\omega)$ be an edge-weighted graph satisfying Setting $\Gamma$]. Then, for all $u,w\in\Gamma$,
@@
where $I$ is the identity matrix in $\mathrm{Mat}_{|E\Gamma|}(\mathbb{C})$ and
@@
In particular, for all $u,w\in\Gamma$ the function $\mathcal{Z}_{\Gamma,u\to w}(s)$ is meromorphic over $\mathbb{C}$.
@@
Fact 5.15 (Matrix Determinant Lemma, cf. [18]).
@@
Let $s\in\mathbb{C}$ be such that $\sum_{n=0}^{\infty}\mathcal{E}(s)^{n}$ converges. Recall that $\sum_{n=0}^{\infty}\mathcal{E}(s)^{n}=(I-\mathcal{E}(s))^{-1}$. By Remark 5.13.1, if $u\in V\Gamma$ then
[…]
@@
[…]
@@
Let $\Gamma$ be a $1$-segment graph with $E\Gamma=\{a,\bar{a}\}$, $c=o(a)$ and $d=t(a)$. Set $\omega(a):=\alpha+1$ and $\omega(\bar{a}):=\beta+1$, where $\alpha,\beta\in\mathbb{Z}_{\geq 1}$ with $\alpha\geq 2$ or $\beta\geq 2$. Set the order $\leq$ on $E\Gamma$ such that $a<\bar{a}$, and identify $\mathbb{C}\llbracket E\Gamma\rrbracket$ with $\mathbb{C}^{2}$, $e_{a}=e_{d}$ with $(1,0)$ and $e_{\bar{a}}=e_{c}=(0,1)$. Then, for every $s\in\mathbb{C}$,
[…]
@@
In particular, let $(G,T)$ be a locally $\infty$-transitive action on a locally finite tree with quotient graph $\Gamma$ and standard edge weight $\omega$ as before. By Remark 5.11.2 and (5.15), we have explicit formulae for $\zeta_{G,G_{v},G_{v}}(s)$ and $\zeta_{G,G_{e},G_{e}}(s)$ for all $v\in VT$ with $G\cdot v=c$ and $e\in ET$ with $G\cdot e=a$. For instance, one may take $G=\mathrm{SL}_{2}(\mathbb{Q}_{p})$ and let $T$ be the Bruhat–Tits tree of $G$ (cf. Example 3.3(ii)). In this case $\alpha=\beta=p$. Let $v\in VT$ be the vertex with $G_{v}=\mathrm{SL}_{2}(\mathbb{Z}_{p})$, and $e\in ET$ be the edge whose pointwise stabiliser is the standard Iwahori subgroup. Then,
@@
Other examples can be obtained from Example 3.12. Note that the formulae before agree with [10, Example 1.7] in case that $G$ is the group of automorphisms of a bi-coloured tree $T$.
@@
[…]
@@ 5.5. Explicit formulae: the (P)-closed case
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram based satisfying Convention 3.1.1. Choose an inversion $\iota$ in $\Delta$ and $c_{0}\in V\Gamma$. Denote by $(T=T(\Delta,\iota,c_{0}),\pi,\mathcal{L})$ the standard $\Delta$-tree associated to $\iota$ and $c_{0}$, let $v_{0}$ be the root of $T$ (cf. Section 3.4.1), and set $G=U(\Delta,\iota,c_{0})$. Assume that the standard edge weight $\omega$ on $\Gamma$ and the standard edge weight $\mathcal{W}$ on $\Delta$ take values in $\mathbb{Z}_{\geq 2}$ and $\mathbb{Z}_{\geq 0}$, respectively. In particular, $T$ has no leaves. Finally, assume that $(G,T)$ has property $(\ast_{k})$ for some $k\geq 1$.
Setting [(P)-cl] guarantees that the series defining $\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)$ converges at some $s\in\mathbb{C}$, for all $t_{1},t_{2}\in T$ (cf. Theorem 5.9). Thanks to Remark 5.0.1, from now on we may focus only on the case in which $t_{1}\in\{v_{0}\}\cup o^{-1}(v_{0})$ while studying $\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)$.
@@
(5.18) $\zeta_{G,G_{v_{0}},G_{t}}(s)=\mathbbm{1}_{\{c_{0}\}}(u)+\hskip-8.5359pt\sum_{a\in o^{-1}(c_{0}),\atop\xi\in\mathcal{P}_{(\Delta,\iota)}(X_{a}\to X_{U})}\hskip-8.5359pt\omega(a)^{-s-1}\mathcal{W}(\xi)^{-s-1}.$
Moreover, for every $e\in o^{-1}(v_{0})$ with $\pi(e)=a$ and $\mathcal{L}(e)=x\in X_{a}$, we have
(5.19) $\zeta_{G,G_{e},G_{t}}(s)=\eta_{a}(u)+\hskip-11.38092pt\sum_{\xi\in\mathcal{P}_{(\Delta,\iota)}(x\to X_{U}),\atop\ell(\xi)\geq 2}\hskip-17.07182pt\mathcal{W}(\xi)^{-s-1}+\hskip-17.07182pt\sum_{y\in X_{c_{0}}\setminus\{x\},\atop\xi\in\mathcal{P}_{(\Delta,\iota)}(y\to X_{U}),\,\ell(\xi)\geq 1}\hskip-28.45274pt\mathcal{W}_{\mathrm{rev}}(x,y)^{-s-1}\mathcal{W}(\xi)^{-s-1},$
@@
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram as in Convention 3.1.1 together with a function $\mathcal{W}\colon X\times X\to\mathbb{Z}_{\geq 0}$ (recall that $X=\bigsqcup_{a\in E\Gamma}X_{a}$). Let $\mathbb{C}\llbracket X\rrbracket$ be the complex vector space of all formal sums $\sum_{x\in X}\gamma_{x}x$, where $\gamma_{x}\in\mathbb{C}$ for every $x\in X$. For every non-empty set $S\subseteq X$, set
$f_{S}:=\sum_{x\in S}x\in\mathbb{C}\llbracket X\rrbracket.$
@@
In what follows, $X$ will be always finite and we implicitly fix a total order on $X$. In this way we can regard $\mathcal{F}(s)$ and the $f_{x}$’s as a $|X|$-dimensional matrix $\mathcal{F}(s)(x,y)]_{x,y\in X}$ and as $|X|$-dimensional row vectors with complex entries, respectively. For all $x,y\in X$, note that $f_{x}\mathcal{F}(s)f_{y}^{t}=\mathcal{F}(s)(x,y)$.
@@
Let $\mathcal{F}(s)$ be as in Definition 5.19 and let $X$ be finite. For every $n\geq 1$, let $\mathcal{F}(s)^{n}$ be the $n$-th power of $\mathcal{F}(s)$, and $\mathcal{F}(s)^{0}$ be the identity matrix in $\mathrm{Mat}_{|X|}(\mathbb{C})$. For $n\geq 0$ and for all $x,y\in X$, we claim that
(5.20) $\mathcal{F}(s)^{n}(x,y)=\sum_{\xi\in\mathcal{P}_{(\Delta,\iota)}(x\to y):\,\ell(\xi)=n+1}\mathcal{W}(\xi)^{-s},$
@@
Let $G=U(\Delta,\iota,c_{0})$ and $T=T(\Delta,\iota,c_{0})$ be as in Setting [(P)-cl]. Let $t\in T$ with $\pi(t)=u$ and let $e\in o^{-1}(v_{0})$ have $\pi(e)=a$ and $\mathcal{L}(e)=x$. Then, for every $r\in\{v_{0},e\}$, we have
@@
$\kappa_{\pi(r)}(u)=\left\{\begin{array}[]{cl}\mathbbm{1}_{\{c_{0}\}}(u)-1,&\text{if }r=v_{0};\\ \mathbbm{1}_{A}(u)-1,&\text{if }r=e.\end{array}\right$.
@@
One proceeds analogously as in the proof of Theorem 5.14. Let $s\in\mathbb{C}$ such that $\sum_{n=0}^{\infty}\mathcal{F}(s+1)^{n}$ converges. By Proposition 5.18 and Remark 5.20.1, we deduce what follows:
$\begin{split}\zeta_{G,G_{v},G_{t}}(s)&=\mathbbm{1}_{\{c_{0}\}}(u)+\sum_{a\in o^{-1}(c_{0})}\omega(a)^{-s-1}f_{X_{a}}\Bigg(\sum_{n=0}^{\infty}\mathcal{F}(s+1)^{n}\Bigg)f_{X_{U}}^{t};\\ \zeta_{G,G_{e},G_{t}}(s)&=\eta_{a}(u)+f_{x}\Bigg(\sum_{n=1}^{\infty}\mathcal{F}(s+1)^{n}\Bigg)f_{X_{U}}^{t}+\\ &\quad+\sum_{y\in X_{c_{0}}\setminus\{x\}}\mathcal{W}_{\mathrm{rev}}(x,y)^{-s-1}f_{y}\Bigg(\sum_{n=0}^{\infty}\mathcal{F}(s+1)^{n}\Bigg)f_{X_{U}}^{t}=\\ =\mathbbm{1}_{A}(u)&+\Big(f_{x}\mathcal{F}(s+1)+\hskip-11.38092pt\sum_{y\in X_{c_{0}}\setminus\{x\}}\hskip-8.5359pt\mathcal{W}_{\mathrm{rev}}(x,y)^{-s-1}f_{y}\Big)\Bigg(\sum_{n=0}^{\infty}\mathcal{F}(s+1)^{n}\Bigg)f_{X_{U}}^{t}.\end{split}$
Since $\sum_{n=0}^{\infty}\mathcal{F}(s+1)^{n}=(I-\mathcal{F}(s+1))^{-1}$, Fact 5.15 yields the claim. ∎